Suppose that Palestinian government raises the price of bread from NIS 4 to NIS 5 and thus



quantity demanded falls from 5,800 per week to 5,500. Calculate total revenue both before



and after the price change. Compute the price elasticity of demand for bread. What can we



tell about the price elasticity of demand for bread?

1
Expert's answer
2021-12-10T14:33:31-0500

Price of bread increases from NIS 4 to NIS 5

The demand decreases from 5800 to 5500

The revenue before "4\u00d75800=23200"

The revenue after

"5\u00d75500=27500"

The elasticity of demand is therefore

"%change in quantity demanded \u00f7 %change in price" percentage change in demand divide by percentage change of price.

The percentage change in quantity demanded is

"\frac{Qd_2 - Qd_1}{Qd_2+Qd_1/2} =\frac {5500-5800}{5500+5800/2}×100 =-5.3%"

Percentage change in price

"\frac{P_2-P_1}{P_2+P_1/{2}}=
\frac{5-4}{5+4/{2}}×100=22.2%"


"PE_d=\frac{-5.3}{22.2}=-0.23"

Thus the price of elasticity of demand for the bread is inelastic.